2.68 × 10 3 moles of a solution containing an ion An+ require 1.61 × 10-3 moles of 〖MnO〗_4^- for the oxidation of An+ to A O_3^- in acidic medium. What is the value of n ?

Grade:11

1 Answers

Navjyot Kalra
9 years ago
Sol. For the oxidation of An+ as : An+ → 〖AO〗_3^- n-factor = 5 – n ⇒ Gram equivalent of An+ = 2.68 × 10-3 (5 - n) Now equating the above gram equivalent with gram equivalent of KMnO4 : 2.68 × 10-3 (5 - n) = 1.61 × 10-3 × 5 ⇒ n = + 2